OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226

OpenStudy (cwrw238):

a22 = a1 +21d plug in the values given to find the common difference d then a31 = a1 + 30d

OpenStudy (anonymous):

so what is it

OpenStudy (anonymous):

-334 -274 -284 -346

OpenStudy (cwrw238):

we r not here just to give you the answer but to show you how the get the answer

OpenStudy (anonymous):

but it needs to be a negative and i dont know how to get the negative

OpenStudy (cwrw238):

a22 = -226 and a1 = 26 so -226 = 26 + 21d 21d = -252 now divide both sides by 21 will give you d

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

okay i got it

OpenStudy (cwrw238):

right now all you do is plug d = -12 into a31 = 26 + 30* -12

OpenStudy (cwrw238):

well i've done it for you ! jsut use your calculator to work it out